# The Construction of Large-scale Structure Catalogs for the Dark Energy Spectroscopic Instrument

### v1

@@ 1 Introduction
The Dark Energy Spectroscopic Instrument (DESI; [1]) was designed to construct three-dimensional maps of the structure of the Universe. It does so by measuring the spectra of pre-selected galaxy and quasar ‘targets’, from which redshifts can measured. Given a cosmological model, the angular celestial coordinates and redshifts can be transformed to comoving Cartesian coordinates. The resulting distribution (or, more commonly, its summary statistics) can be compared to that predicted by the cosmological model. In this manner, cosmological constraints obtained with DESI, and the results from Data Release 1 (DR1;[2]) are presented in [3, 4, 5, 6, 7, 8].
@@
$\bullet$ The final pieces of the LSS pipeline are described in Section 7, where we describe the production of the ‘clustering’ catalogs that are meant to be input to DESI clustering measurement pipelines. This includes cutting the data to secure redshifts within a desired range, assigning redshifts to the random catalogs, and calculating weights that are meant to optimally account for changes in the observed density, e.g., as a function of redshift and the number of overlapping tiles.
We provide a summary and concluding remarks in Section 8. The final products are meant to easily enable the calculation of clustering statistics, and such 2-point measurements obtained from the DR1 version of DESI LSS catalogs are presented in [13]. For convenience, a glossary defining relevant DESI terms, including those defined during the LSS pipeline process, is presented in an appendix.
@@ 3.1 Collecting Reachable Targets
We use the DESI fiberassign [25] software [^8] https://github.com/desihub/fiberassign to define the DESI footprint and collect the information on all targets within it. The properties of all cases in which a DESI target is reachable to a given fiber positioner are determined for every tile; each case represents a ‘potential assignment’. Collecting this information is required to determine completeness statistics and correct for their variations properly. Each fiber positioner is expected to be able to reach many DESI targets. All of these will be potential assignments, but only one will receive the actual assignment, which is given to the target that has the highest PRIORITY (and then SUBPRIORITY if multiple have the same PRIORITY). The potential assignments information is recorded within the same file that designates the final (actual) fiber positioner assignments999within the POTENTIAL_ASSIGNMENTS HDU, for the files described here: https://desidatamodel.readthedocs.io/en/latest/DESI_TARGET/fiberassign/tiles/TILES_VERSION/TILEXX/uncompressed-fiberassign-TILEID.html.
@@ 4 Target Catalogs in DESI Footprint
In the previous section, we detailed how the information associated with each case in which a target was reachable is gathered and stored in combined target and spectroscopic information catalogs, typically resulting many entries per TARGETID. The next step in the DESI LSS pipeline is to produce the ‘full’ catalogs that contain entries for all reachable targets, whether or not they were observed, but are cut to unique TARGETID. These catalogs are produced per supported tracer type (described in 3). We describe the process of creating these catalogs and the stages they go through in this section. Weights to account for fluctuations in the selection function are eventually added to these catalogs but will described in Section 5.
@@ 4.1 Full catalogs, Without Vetos
•  $T_{\rm fa}$: The particular combination of fiber and tile (but not necessarily target) was assigned and observed; i.e., for the LRG sample, this quantity will be True if any LRG was assigned an observed on the given fiber and tile, but, given that we have already cut to LRG targets at this point, it will be False if any other type of target was assigned to the particular combination of fiber and tile (the subscript fa stands for ‘fiber assigned’).
@@ 4.2 Applying Vetos
Fig. 3 shows the angular effect of veto masks applied to the DR1 LRG data and random samples, within a small area of the footprint. Unmasked data can be seen with black points. The yellow points are areas vetoed by due to the priority of what was assigned blocking the possibility of any LRG assignments. This primarily happens in areas that have only been covered by one tile; one can observe that significant amounts of such areas are yellow. The blue points denote data where some aspect of the DESI instrument was flagged as poorly performing and thus denoted ‘bad hardware’ (i.e., $H_{\rm gd}=\texttt{False}$, as defined above). In the lower left-hand corner of each of the upper panels, one can observe that an entire petal was so flagged. Looking again at Fig. 2, one can see no data in the location for DR1 and slightly lower coverage than average as of March 20th, 2024. This is because data with $H_{\rm gd}=\texttt{False}$ does not contribute to our coverage determination. At other locations, smaller-sized patterns are removed; these can correspond to, e.g., amplifier boundaries on the CCDs that read out the information from the spectrograph or groups of 50 fibers that failed to move properly due to canbus communication errors. Similar to the priority mask, such areas exist primarily where only one tile has covered the area. For the vast majority of the combination of exposures and fibers the DESI hardware performs well and failure is thus especially unlikely for any area covered more than once. In total, in DR1, only 3% and 2% footprint is vetoed due to poorly performing hardware, for dark- and bright-time respectively, and these numbers are expected to decrease in future data releases as the coverage increases (when the same area is covered again, it is unlikely the hardware will again be poorly performing, especially given it will be a different area in the focal plane). All of this information is so tracked and incorporated into the LSS catalog footprint.
@@ 5.2 Density Variations due to Imaging
  • The neural net method sysnet232323 https://github.com/mehdirezaie/sysnetdev, most recently applied to [30] and denoted WEIGHT_SN in the catalogs;
@@ 7 Clustering catalogs
The DESI LSS ‘clustering’ catalogs cut to good spectroscopic observations and the redshift range intended for clustering analysis. They only include the columns required to calculate clustering statistics.
@@ 7.1 Reducing Data to what is used for Clustering
All weights are kept in the ‘clustering’ catalogs to make it easy to test the impact of ignoring a given weight (e.g., one would divide the $w_{\rm tot}$ column by $w_{\rm zfail}$ when running clustering measurements to these the effect of ignoring redshift failure weights). Similarly, when multiple versions of the regressions against imaging properties have been run, we provide each of the individual weights for testing. Additionally, we keep the number of overlapping tiles (NTILE) to allow tests of clustering as a function of coverage. The comprehensive list with definitions will be updated with each data release.262626 https://desidatamodel.readthedocs.io/en/latest/DESI_ROOT/vac/RELEASE/lss/VERSION/LSScats/clustering/clustering_dat.html Any desired columns not provided can be obtained via a TARGETID match to the ‘full’ catalogs (or any other DESI catalogs).
@@ 8 Summary
The steps in the DESI LSS catalog pipeline can be summarized as:

### v2

@@ 1 Introduction
The Dark Energy Spectroscopic Instrument (DESI; [1]) was designed to construct three-dimensional maps of the structure of the Universe. It does so by measuring the spectra of pre-selected galaxy and quasar ‘targets’, from which redshifts can be measured. Given a cosmological model, the angular celestial coordinates and redshifts can be transformed to comoving Cartesian coordinates. The resulting distribution (or, more commonly, its summary statistics) can be compared to that predicted by the cosmological model. In this manner, cosmological constraints obtained with DESI, and the results from Data Release 1 (DR1;[2]) are presented in [3, 4, 5, 6, 7, 8].
@@
$\bullet$ The final pieces of the LSS pipeline are described in Section 7, where we describe the production of the ‘clustering’ catalogs that are meant to be input to DESI clustering measurement pipelines. This includes cutting the data to secure redshifts within a desired range, assigning redshifts to the random catalogs, and calculating weights meant to optimally account for changes in the observed density, e.g., as a function of redshift and the number of overlapping tiles.
We provide a summary and concluding remarks in Section 8.
The clustering catalog products are meant to enable the simple calculation of clustering statistics. The detailed information on the pipeline process presented in this work and its various assumptions allows biases in the resulting clustering statistics to be predicted and modeled. However, the details of such modeling are presented elsewhere. The specific choices regarding the pipeline applied to the DR1 data and the resulting 2-point clustering measurements [^2] This includes various window functions required to translate between true and observed redshift-space clustering measurements. are presented in [13]. The ability to extract unbiased cosmological information from DESI LSS catalogs obtained from realistic simulations is the subject of many DESI DR1 studies, including [14, 15, 16].
For convenience, a glossary defining relevant DESI terms, including those defined during the LSS pipeline process, is presented in an appendix.
@@ 3.1 Collecting Reachable Targets
We use the DESI fiberassign [28] software [^9] https://github.com/desihub/fiberassign to define the DESI footprint and collect the information on all targets within it. The properties of all cases in which a DESI target is reachable to a given fiber positioner are determined for every tile; each case represents a ‘potential assignment’. Collecting this information is required to determine completeness statistics and correct for their variations properly. Each fiber positioner is expected to be able to reach many DESI targets. All of these will be potential assignments, but only one will receive the actual assignment, which is given to the target that has the highest PRIORITY (and then SUBPRIORITY if multiple have the same PRIORITY). The potential assignments information is recorded within the same file that designates the final (actual) fiber positioner assignments101010within the POTENTIAL_ASSIGNMENTS HDU, for the files described here: https://desidatamodel.readthedocs.io/en/latest/DESI_TARGET/fiberassign/tiles/TILES_VERSION/TILEXX/uncompressed-fiberassign-TILEID.html.
@@ 4 Target Catalogs in DESI Footprint
In the previous section, we detailed how the information associated with each case in which a target was reachable is gathered and stored in combined target and spectroscopic information catalogs, typically resulting in many entries per TARGETID. The next step in the DESI LSS pipeline is to produce the ‘full’ catalogs that contain entries for all reachable targets, whether or not they were observed, but are cut to unique TARGETID. These catalogs are produced per supported tracer type (described in 3). We describe the process of creating these catalogs and the stages they go through in this section. Weights to account for fluctuations in the selection function are eventually added to these catalogs but will described in Section 5.
@@ 4.1 Full Catalogs, Without Vetos
•  $T_{\rm fa}$: The particular combination of fiber and tile (but not necessarily target) was assigned and observed; i.e., for the LRG sample, this quantity will be True if any LRG was assigned and observed on the given fiber and tile, but, given that we have already cut to LRG targets at this point, it will be False if any other type of target was assigned to the particular combination of fiber and tile (the subscript fa stands for ‘fiber assigned’).
@@ 4.2 Applying Vetos
Fig. 3 shows the angular effect of veto masks applied to the DR1 LRG data and random samples, within a small area of the footprint. Unmasked data can be seen with black points. The yellow points are areas vetoed due to the priority of what was assigned blocking the possibility of any LRG assignments. This primarily happens in areas that have only been covered by one tile; one can observe that significant amounts of such areas are yellow. The blue points denote data where some aspect of the DESI instrument was flagged as poorly performing and thus denoted ‘bad hardware’ (i.e., $H_{\rm gd}=\texttt{False}$, as defined above). In the lower left-hand corner of each of the upper panels, one can observe that an entire petal was so flagged. Looking again at Fig. 2, one can see no data in the location for DR1 and slightly lower coverage than average as of March 20th, 2024. This is because data with $H_{\rm gd}=\texttt{False}$ does not contribute to our coverage determination. At other locations, smaller-sized patterns are removed; these can correspond to, e.g., amplifier boundaries on the CCDs that read out the information from the spectrograph or groups of 50 fibers that failed to move properly due to canbus communication errors. Similar to the priority mask, such areas exist primarily where only one tile has covered the area. For the vast majority of the combination of exposures and fibers the DESI hardware performs well and failure is thus especially unlikely for any area covered more than once. In total, in DR1, only 3% and 2% footprint is vetoed due to poorly performing hardware, for dark- and bright-time respectively, and these numbers are expected to decrease in future data releases as the coverage increases (when the same area is covered again, it is unlikely the hardware will again be poorly performing, especially given it will be a different area in the focal plane). All of this information is so tracked and incorporated into the LSS catalog footprint.
@@ 5.2 Density Variations due to Imaging
  • The neural net method sysnet242424 https://github.com/mehdirezaie/sysnetdev, most recently applied to [33] and denoted WEIGHT_SN in the catalogs;
@@ 7 Clustering Catalogs
The DESI LSS ‘clustering’ catalogs are cut to good spectroscopic observations and the redshift range intended for clustering analysis. They only include the columns required to calculate clustering statistics.
@@ 7.1 Reducing Data to what is used for Clustering
All weights are kept in the ‘clustering’ catalogs to make it easy to test the impact of ignoring a given weight (e.g., one would divide the $w_{\rm tot}$ column by $w_{\rm zfail}$ when running clustering measurements to these the effect of ignoring redshift failure weights). Similarly, when multiple versions of the regressions against imaging properties have been run, we provide each of the individual weights for testing. Additionally, we keep the number of overlapping tiles (NTILE) to allow tests of clustering as a function of coverage. The comprehensive list with definitions will be updated with each data release.272727 https://desidatamodel.readthedocs.io/en/latest/DESI_ROOT/vac/RELEASE/lss/VERSION/LSScats/clustering/clustering_dat.html Any desired columns not provided can be obtained via a TARGETID match to the ‘full’ catalogs (or any other DESI catalogs).
@@ 8 Summary
Figure 4: A flowchart detailing the DESI LSS catalog pipeline construction steps. The full context, including key definitions, can be found within the labeled sections of this paper.
A flowchart detailing the steps in the DESI LSS catalog pipeline is shown in Fig. 4. We summarize these steps as: